OpenStudy (anonymous):

A plane fare including 8% in taxes and fees, is $280.80. How much are the taxes and fees?

OpenStudy (anonymous):

@dmezzullo

OpenStudy (dmezzullo):

we did this yest divide 280.80 by 8

OpenStudy (anonymous):

OHH RIGHT! sorry! xD It looked different for some reason.

OpenStudy (dmezzullo):

its cool. tag me if u got more

OpenStudy (anonymous):

What?? no don't divide by 8! x + x*0.08 = 280.80 Solve for x

OpenStudy (anonymous):

x will the original cost. If you multiply your value of x by 0.08, that will be the taxes and fees

OpenStudy (anonymous):

You know what the funny thing is? in mu book it says the answer is $20.80. I'm trying to figure out how they go that. :0

OpenStudy (anonymous):

Thank you @dmezzullo

OpenStudy (anonymous):

I will :)

OpenStudy (dmezzullo):

Np.

OpenStudy (anonymous):

Dmezzullo is completely wrong!!!! Look: x + x*0.08 = 280.80 1.08x = 280.80 x=260 So the original cost was 260$ Now multiply 260 by 0.08 260*0.08 = 20.8$

OpenStudy (anonymous):

@dmezzullo don't help people if you have no idea what you're doing, you're misleading people

OpenStudy (anonymous):

ok Thank you this helps. :) I they meant too.:)

OpenStudy (dmezzullo):

How am i misleading ppl explain tht to me.

OpenStudy (anonymous):

Dividing by 8 is completely useless, has nothing to do with the question or the answer and gives you an incorrect answer.

OpenStudy (anonymous):

oops! sorry I meant to say i dont think he meant too.

OpenStudy (anonymous):

I recommend you unfan dmezzullo, @GodBlessYou, he's a fraud! :P

OpenStudy (dmezzullo):

How am i a fraud? @BluFoot

OpenStudy (anonymous):

I just told you!! See my previous post about how dividing by 8 is useless

OpenStudy (dmezzullo):

EXPLAIN TO ME WHAT I DID WRONG! @BluFoot SINCE U KNOW EVERYTHING MR KNOW IT ALL.

OpenStudy (anonymous):

LOl! I think they just made a mistake. Its not worth aruging over. :D

OpenStudy (dmezzullo):

Don't u got some explaining to do @BluFoot

OpenStudy (anonymous):

@dmezzullo OK 280.80$ is including sales and taxes. Which means that the original price + the taxes = 280.80 so if we call the original price x then x + x*8/100 = 280.80 Solve for x and you get the original price. Multiply that value by 8, divide by 100 and you get the taxes

OpenStudy (anonymous):

I dont think your a fraud @dmezzullo :) Everyone miscalculates sometime. ;)

OpenStudy (anonymous):

Dividing 280.80 by 8 gives you 35.1, but the answer is 20.8. Clearly does not work

OpenStudy (anonymous):

It's not a miscalculation, it's a completely incorrect method of approaching the problem

OpenStudy (dmezzullo):

Or its called im a human that makes mistakes.... Not a "Mr. Know It All Like U."

OpenStudy (dmezzullo):

@SWAG wat would u say bout this? Tiny mistake?

OpenStudy (anonymous):

Hey! SWAG just changed his picture! lol! :)

OpenStudy (swag):

@GodBlessYou Yes i did & @BluFoot No need to be rude to @dmezzullo

OpenStudy (anonymous):

Can we be done with this ong drawn out discussion? lol:D Thank you for your guys help! I am going to try to write this our and figure it out. :)

OpenStudy (anonymous):

i mean long discussion sorry.